OpenStudy (anonymous):

Set up a linear system consisting of two equations. Assume you will talk for a minimum of 600 minutes. The first equation would be for the Talks-A-Lot Company. The total cost, y, equals the base fee plus cost per minute times the number of minutes exceeding 600 minutes. The second equation would be set up just like the first, only you need to use the information for the Chat-Away Company.

OpenStudy (anonymous):

Solve the linear system using the substitution method. Please make sure to solve for both x and y. Show all work.

OpenStudy (anonymous):

Answer the questions, using complete sentences. How many minutes would you have to talk over and above the 600 minutes for the cost to be the same with both companies? What would be the cost when the minutes are the same?

OpenStudy (anonymous):

@leedomathgeek

OpenStudy (anonymous):

Let "a" denotes the fees for each minute exceeding the 600 minutes and let "f1" be the fees of the monthly plane for the 600 minutes. If "x" represents the number of extra minutes then the total cost would be y=ax+f1

OpenStudy (anonymous):

okay so now what

OpenStudy (anonymous):

For the second company you have the same equation but different fees parameters, say y=bx+f2

OpenStudy (anonymous):

so y = 600(x) + f2?

OpenStudy (anonymous):

"a" represents the fees for an extra minute beyond 600 minutes

OpenStudy (anonymous):

For example if the fee is a=$0.25, and the monthly plan costs $35 per month, then y=0.25*x+35

OpenStudy (anonymous):

This will give you the total cost for talking beyond 600 minutes (which are included in your $35 plan)

OpenStudy (anonymous):

If x=0 (no extra minutes), then the total cost per month is just $35 (your monthly plan)

OpenStudy (anonymous):

so can u help me come up with a linear equation

OpenStudy (anonymous):

You need to specify the monthly plan, and the extra charge rate? Are they given in the question?

OpenStudy (anonymous):

You just purchased a cellular phone and are trying to determine to which cell phone company you will give your business. When you contacted the Talks-A-Lot Company, they were offering a monthly plan of $40 for 600 minutes and $0.35 for each minute exceeding the 600 minutes. In the Sunday paper you see an ad for the Chat-Away Company, which offers a monthly plan of $50 for 600 minutes and $0.10 for each minute exceeding the 600 minutes. How many minutes would you have to talk over and above the 600 minutes for the cost to be the same with both companies? What would be the equal cost?

OpenStudy (anonymous):

Exactly...so for the first company you have y=0.35(x)+40 and for the second one you have y=0.1(x)+50

OpenStudy (anonymous):

y=0.35(x)+40 and y=0.1(x)+50 is answer to part 1 correct

OpenStudy (anonymous):

For the cost to be the same with both companies your need to let y1=y2 0.35x+40=0.1x+50 0.25x=10 x=10/.25=40 minutes

OpenStudy (anonymous):

And the equal cost would be y=0.1*(40)+50=$54

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait wait..

OpenStudy (anonymous):

y=0.35(x)+40 and y=0.1(x)+50 is the linear equations for part 1 correct or are we stilll trying to do linear equations

OpenStudy (anonymous):

so now i gotta solve both of those linear equations by substitution correct by the directions

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Exactly, the first equation is for the first company and second equation is for the second company

OpenStudy (anonymous):

okay so lets work one by one.

OpenStudy (anonymous):

Those are two linear equations...

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

okay so lets work the first one out using substitution method

OpenStudy (anonymous):

can u help me work through it

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

y1=0.35x+40 y2=0.1x+50

OpenStudy (anonymous):

we need to find the value of x that makes y1 and y2 equal

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let y1=y2=a (a is the value that we need to find)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

In this case, we have from the first equation a=0.35x+40 ---> solve for x 0.35x=40-a x=(40-a)/0.35

OpenStudy (anonymous):

Do you have any question so far?

OpenStudy (anonymous):

no

OpenStudy (goformit100):

it's correct

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

Then, substitute x that was found previously in the second equation we get a=0.1[(a-40)/0.35]+50 a=(10/35)(a-40)+50 (multiply both sides by 35) 35a=10(a-40)+50*35=-400+1750+10a 25a=2150 a=1350/25 =$54

OpenStudy (anonymous):

How many minutes would you have to talk over and above the 600 minutes for the cost to be the same with both companies? The answer is $54

OpenStudy (anonymous):

I hope you got it now?

OpenStudy (anonymous):

yes i got it

OpenStudy (anonymous):

so that solves for part 2 correct answer is $54

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and for part 3 part a would be 54 minutes corect

OpenStudy (anonymous):

This is the equal cost

OpenStudy (anonymous):

so how part part 3 section a question

OpenStudy (anonymous):

Answer the questions, using complete sentences. How many minutes would you have to talk over and above the 600 minutes for the cost to be the same with both companies? What would be the cost when the minutes are the same?

OpenStudy (anonymous):

The minutes is x=(a-40)/0.35=(54-40)/0.35=40 minutes

OpenStudy (anonymous):

so u would have to talk over 40 minutes for the cost to be the same with both companies?

OpenStudy (anonymous):

It is my bad ..I wrote before 54 ....

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

ok now what about part b

OpenStudy (anonymous):

What would be the cost when the minutes are the same?

OpenStudy (anonymous):

If the total cost is y=0.35x+40 --> then substitute x=40minutes you get.....y=$54

OpenStudy (anonymous):

so 54 dollers would be the cost when the minutes are the same

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you may plug x=40 minutes into both equations and you will get the same answer

OpenStudy (anonymous):

If you plan to talk for 1000 minutes, which company should you hire? Please show your total cost for both companies to prove your answer.

OpenStudy (anonymous):

OK then, in this case, your extra minutes are x=1000-600=400 minutes. For company 1: the total cost would be y=0.35*400+40=$180 For company 2: the total cost would be y=0.1*400+50=$90 In this case, I would hire the second company

OpenStudy (anonymous):

the second company is Chat-Away Company

OpenStudy (anonymous):

Yup...Chat-Away

OpenStudy (anonymous):

thanks for the help..

OpenStudy (anonymous):

no problem....good luck

OpenStudy (anonymous):

can u help me with something else

OpenStudy (anonymous):

Sure..Let me see the question first and then I will let you know...I have few minutes to spare